satellite73 (satellite73):

ok i am still stuck. expand \[\frac{z^2-1}{(z^2+1)^2}\] at \[z_0=i\]

OpenStudy (anonymous):

jamesj suggesting rewriting as \[\frac{1}{(z-i)^2}\frac{z^2-1}{(z+i)^2}\] but the second part is killing me

OpenStudy (anonymous):

Satellite I didn't know u ever needed help

OpenStudy (anonymous):

for some reason this laurant expansion is kicking my butt. i know the answer, but i don't know a good way to get it

OpenStudy (anonymous):

I can't even figure out how jamesj got that!!!!!!

OpenStudy (anonymous):

denominator is \[(z^2+1)^2=(z+i)(z-i)^2=(z-i)^2(z+i)^2\] is how,but the expansion is eluding me

OpenStudy (matt101):

That isn't the right factorization of the denominator. You used difference of squares, but a sum of squares cannot be factored.

OpenStudy (anonymous):

@jamesj i tried what you suggested, but it is not clear to me how dividing \[\frac{z^2-1}{(z+i)^2}\] get me the expansion. i just get a quotient of 1 and a remainder of something, i forget what

OpenStudy (mathteacher1729):

Satellite -- I'm at work right now so I don't have much time, but if you haven't already, you should check out this free online set of notes about complex analysis. It's rigorous and readable. The guy taught it for 20 years, so he knows his stuff : http://rutherglen.science.mq.edu.au/wchen/lnicafolder/lnica.html Hope this helps! P.S. He has other notes as well: http://rutherglen.science.mq.edu.au/wchen/ln.html Cheers!

OpenStudy (anonymous):

thanks for the site

OpenStudy (anonymous):

@mathteacher, thanks i will look. but there is something in this computation i am missing. i have to run but will be back. thanks

OpenStudy (jamesj):

\( (z+i)^2 = z^2 + 2iz - 1 \) hence \[ \frac{z^2 - 1}{(z+i)^2} = 1 - \frac{2iz}{(z+i)^2} \] and \[ \frac{2iz}{(z+i)^2} = (z-i).\frac{2i}{(z+i)^2} + \frac{-2}{(z+i)^2} \] Now the Laurent expansions of \( 1/(z+i)^2 \) are easy to evaluate. Find them, multiply by the appropriate constants, shift them down given the powers of (z-i), put it all back together.

OpenStudy (anonymous):

ok now i think i got it finally. sorry to be so slow.

OpenStudy (anonymous):

i think i was looking for it to be somewhat easier and that got me stuck

OpenStudy (jamesj):

This one isn't so bad. Complex analysis is fun and very powerful; hopefully you'll finish your course with that sense.